OpenStudy (anonymous):

Assume x and y are functions of t. Evaluate dy/dt for the following: 4x^3-9xy^2+y=-80; dx/dt= 4, x=-3, y=1

OpenStudy (anonymous):

y=-82

OpenStudy (anonymous):

y+82=0

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

It's easy just use combine like terms than you evaluate

OpenStudy (anonymous):

ah, my book says we are supposed to first differentiate and then plug in the values

OpenStudy (anonymous):

\[4x^3-9xy^2+y=-80\] \[12x^2x'-9(x'y^2+2xyy')+y'=0\] solve for y'

OpenStudy (anonymous):

K, ill do that rly quick

OpenStudy (anonymous):

or replace x by -3, y by 1, x' by 4 and then solve for y'

OpenStudy (anonymous):

ah that might be easier

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yes probably easier with the numbers